# Multi-tracer beyond linear theory

### v1

@@ 3.3 The multi-tracer information content
which represents the average squared value of the non-diagonal elements, with $N$ being the total number of parameters. Lower values indicate a more orthogonal basis for the parameter $i$. We display this score for $h$ in the top right panel of Fig. 5, showing results for three cases: linear modeling with a split in $b_{1}$, nonlinear modeling with a split in $b_{1}$ and nonlinear modeling with a split in $b_{\rm all}$. For the linear case, we find a small improvement of MT when breaking degeneracies compared to the ST scenario, but we could only find a very mild dependence with $\Delta b_{1}$. In the non-linear modeling, MT exhibits a clear degeneracy break compared to ST, but we also find strong dependence in $\Delta b$. This indicates that a large difference in the bias leads to a more orthogonal basis for the cosmological parameters. The dependence on $\Delta b_{\rm all}$ is even stronger than on $\Delta b_{1}$, indicating that using all bias parameters together may lead to greater degeneracy breaking, particularly coming from the non-linear terms.
@@ 4.1 More than two tracers
Figure 6: Relative errors in $\omega_{\rm cdm}$, $h$ and $A_{s}$ as a function of the number of tracers $N_{\rm Tracers}$ and for different $\bar{n}^{T}$ values. We fix $b_{\rm all}=0.6$. For the solid lines, we include the cross-stochastic terms and we exclude them for the dashed lines.
We display as solid lines in Fig. 6 the evolution of the Fisher estimated error in $\omega_{\rm cdm}$, $h$ and $A_{s}$ as a function of the number of tracers, having $b_{\rm all}=0.6$ fixed. For the tracer split with more tracers, we generalize Eq. (2.27) using evenly spaced MT bias parameters
@@ 4.2 The dependence on the total and relative number density
Figure 7: In the left panel, the relative error in the cosmological parameters as a function of the total number density $\bar{n}^{T}$. Dashed lines correspond to linear theory while solid lines represent the EFT non-linear theory. For the EFT MT analysis, we consider splits in $b_{1}$ and $b_{\rm all}$. In the right panel, the same results normalized by the corresponding ST scenarios. We fix $b_{1}=0.6$ or $b_{\rm all}=0.6$.
@@
Figure 8: Error in the cosmological parameters normalized by the ST error as a function of the fraction of the total number density in tracer $t_{1}$ for a non-balanced split, defined as $f_{t_{1}}=\bar{n}^{t_{1}}/\bar{n}^{T}$. Different colors correspond to different values of $\bar{n}^{T}$, in units of $h^{3}/$Mpc3. As before, we fix $b_{\rm all}=0.6$.
When performing a MT split, we also have the freedom to chose the number densities of the final tracers with respect to the initial sample. In [45, 46] we have only consider a balanced split, in which each subtracer has the same number density. Here we extend this analysis by studying the MT performance as a function of the tracer $t_{1}$ fraction, defined in Eq. (3.2). In this work, we keep the MT relative bias difference $\Delta b$ constant with $f_{t_{1}}$. However, it is important to note that in realistic scenarios where the two samples are split based on a tracer feature, $\Delta b$ is expected to change with $f_{t_{1}}$ (see Sec. 3.2). Since this dependence heavily depends on the sample considered, we assume for simplicity a constant bias difference. We display in Fig. 8 the error as a function of $f_{t_{1}}$ for different values of $\bar{n}^{T}$ normalized by the ST error. Note that the case $f_{t_{1}}\to 0$ or $f_{t_{1}}\to 1$, for which the full density is in one of the samples, does not reproduce the ST case for the same fiducial bias parameters Eq. (2.26), but rather corresponds to a different set of bias parameters shifted by $\Delta b$. We see a typical $U$-shape curve with a relatively flat region for $0.1\lesssim f_{t_{1}}\lesssim 0.9$, which is especially present as we increase $\bar{n}^{T}$. It indicates that, despite the optimal value normally being close to the balanced case $f_{t_{1}}=0.5$, the MT improvements are also present with some unbalancing between both tracer densities. Notice that, in general, it is easier to find subsamples with distinct bias parameters if allowing for their number densities to not be exactly the same. One can, for instance, find the 10 or $20\%$ galaxies with lower or higher values for $b_{\mathcal{G}_{2}}$ and already benefit from the MT information gains.
@@ 5 FoG and different $k_{\rm max}$
[…]
@@
Table 2: Different $k_{\rm max}$ values (in units of $h/$Mpc) used for the auto, cross and single tracer spectra based on the zero-crossing of the quadrupole.
@@ 6 Forecasts
summing over multiple redshift bins $z_{i}$ and anchoring the power spectra at $k=0.1\,h/$Mpc. In addition, we compute the effective number density $\overline{n}_{\text{eff}}$ and the effective redshift $z_{\text{eff}}$ by averaging $\overline{n}(z_{i})$ and $z_{i}$ weighted by the number of tracers $N_{i}=\overline{n}(z_{i})V(z_{i})$ in the corresponding redshift bin. For the linear bias, we use $b_{1}^{\rm eff}(z_{\rm eff})=0.9+0.4z_{\rm eff}$, which has been used in other works [8, 95, 96] and is based on [97]. For simplicity, we fix the other non-linear bias parameters to Eq. (2.26). We summarize in Tab. 3 the surveys and their specifications used in this section.
@@
The forecast results are presented in Fig. 10 as a function of $\Delta b_{\rm all}$. The solid lines represent the most conservative scenario in which we include all the terms, while the dashed lines represent the most optimistic case in which the cross-stochastic component is neglected. We find a factor 2 improvement in $A_{s}$ across all surveys. We also find significant improvement in $h$ and $\omega_{\rm cdm}$, but somewhat smaller for PFS and MegaMapper, which we checked to be due the high $z$ considered for those surveys. The non-linear corrections are very small at high redshift at $k=0.15\,h/$Mpc, and typically one can reach higher values of $k_{\rm max}$ at high $z$. A more comprehensive study which considers the different redshift bins of each survey and models their respective $k_{\rm max}$ values as a function of redshift, is left for future work. Finally, when the cross-stochastic component is neglected, we find even stronger MT constraints, highlighting the importance of adding simulation-based priors on these parameters.
@@ 7 Conclusions
In this article, we have extended the results of [45, 46] carrying a systematic study of MT when considering the non-linear modeling of galaxy clustering. We employ a series of Fisher studies to investigate different scenarios, including various survey configurations. Within the large-scale bias expansion, any tracer split based on a specific sample feature can be mapped into a direct split in the bias parameters. Accordingly, we display our results as a function of the difference $\Delta b$ in the bias coefficients and show that the MT Fisher information reproduces the ST scenario when taking $\Delta b\to 0$. Further, we discuss how likely it is to identify two halo samples with differing non-linear bias parameters. While such differences are hard to find in a vanilla scenario, the presence of assembly bias and the use of realistic galaxy samples can lead to $\Delta b_{\mathcal{G}_{2}}\sim 1$ [63]. This paper paves the way toward identifying a MT bias split that does not necessarily rely on finding two samples with very different linear bias coefficients.
@@
Next, we study the importance of cross-stochasticity for MT. We show that neglecting this term is vital for MT within linear theory, while this appears not to be the case when considering the non-linear bias modeling. Furthermore, we show for the first time, using non-linear modeling, that a split into two tracers seems to maximize the information gain, at least when restricted to a power spectra analysis. We have also found that, when considering the non-linear operators, the MT gains are present at very realistic number density values, in contrast with the findings on linear theory that demands very high tracer densities. We addressed the question of what are the optimal number densities between the two subtracers, finding substantial gains even when considering a very unbalanced split, i.e. when one the tracers is more dense than the other. This result makes the task of finding two tracers with two different set of biases simpler, since we can limit the MT searches to find a small in-homogeneous subsample. Finally, in Sec. 5 we consider the possibility of using different values of $k_{\rm max}$ for the subtracers depending on the strength of the FoG effect for each. We show how this approach can substantially enhance the information extracted from the bluer sample.
@@ Acknowledgments
We thank Raul Abramo, David Alonso, Anna Cremaschi, Boryana Hadzhiyska, Steffen Hagstotz, Eiichiro Komatsu, Fiona McCarthy, Thiago Mergulhão, Srinivasan Sankarshana, Barbara Sartoris, Fabian Schmidt, Blake Sherwin, Beatriz Tucci and Matteo Zennaro for useful discussions. We thank Alex Barreira, Mathias Garny and Rodrigo Voivodic for valuable comments in the draft.
@@ Appendix A Stability of Fisher derivatives
Figure 11: In the left panel, the relative error $\sigma_{h}/h^{\rm fid}$, $\sigma_{\omega_{\rm cdm}}/\omega_{\rm cdm}^{\rm fid}$ and $\sigma_{A_{s}}/A_{s}^{\rm fid}$ as a function of the relative step size $\delta h$, $\delta\omega_{\rm cdm}$ and $\delta A_{s}$ used in the (second-order) finite difference method. A small step size is consistent with noise while a large step size leads to the breakdown of the finite difference method. We adopt $\delta h/h^{\rm fid}=\delta\omega_{\rm cdm}/\omega_{\rm cdm}^{\rm fid}=\delta A_{s}/A_{s}^{\rm fid}=0.04$, where the errors remain relatively stable. In the right panel, the ratio between MT and ST error bars as a function of the fiducial bias parameters chosen in Eq. (2.26). For the MT results, we consider $\Delta b=0.3$ (dashed) and $\Delta b=0.6$ (solid) splits.
The derivatives $\partial_{\theta_{a}}P^{t_{i}t_{j}}(k_{\alpha})$ in the Fisher analysis of Eq. (2.18) can be computed analytically with respect to the bias and stochastic parameters but have to be computed numerically for $h$, $\omega_{\rm cdm}$ and $A_{s}$. To improve the stability of the numerical derivatives, we use a second-order finite difference method and compute the log-derivative $P^{t_{i}t_{j}}(k_{\alpha})\partial_{\theta_{a}}\ln P^{t_{i}t_{j}}(k_{\alpha})$. We show in the left panel of Fig. 11 the stability of the Fisher-extracted relative parameters errors $\sigma_{h}$, $\sigma_{\omega_{\rm cdm}}$ and $\sigma_{A_{s}}$ as a function of the step sizes $\delta h$, $\delta\omega_{\rm cdm}$ and $\delta A_{s}$ used for the finite difference evaluation. We observe that choosing a step size that is too small results in numerical noise, while large step sizes lead to the breakdown of the finite difference method. Intermediate values lead to relatively stable results. Based on this, we adopt the relative step size $\delta h/h^{\rm fid}=\delta\omega_{\rm cdm}/\omega_{\rm cdm}^{\rm fid}=\delta A_{s}/A_{s}^{\rm fid}=0.04$, normalized by their fiducial Planck 18 values.

### v2

@@ 3.3 The multi-tracer information content
which represents the average squared value of the non-diagonal elements, with $N$ being the total number of parameters. Lower values indicate a more orthogonal basis for the parameter $i$. We emphasize that a lower score does not necessarily implies better or worst constraining power, which is ultimately determined by the Fisher matrix $F$, but it can serve as a diagnostic to check whether degeneracy breaking can explain differences in the Fisher information. We display this score for $\omega_{\rm cdm},h$ and $A_{s}$ in the top, middle and bottom panels of Fig. 6, showing results for three cases: linear modeling with a split in $b_{1}$ (left) and nonlinear modeling with a split in $b_{1}$ or $b_{\rm all}$ (right). The dotted lines are obtained discarding the cross-stochastic contribution. We see that removing the cross-stochasticity dramatically improves the degeneracy score in both linear and non-linear cases (with the exception of $S_{h}$ in the linear regime, which remains roughly unchanged). This indicates that in the absence of cross-stochasticity, the cross-spectra information substantially enhances the MT performance. As previously mentioned, adding motivated priors on cross-stochasticity may therefore help break degeneracies between parameters in MT.
In the linear case (left), we find almost no dependence on $\Delta b_{1}$ for all parameters, indicating that a large $b_{1}$ difference between tracers induces almost no degeneracy breaking. On the other hand, removing the cross-stochastic parameter significantly decreases the degeneracy score compared to ST for $A_{s}$ and $\omega_{\rm cdm}$, suggesting that the cross-spectra information in this case serves to break degeneracies and can explain the MT gains.
In the non-linear modeling (right), removing the cross-stochasticity leads to major gains in degeneracy breaking, in agreement with the results presented in Fig. 5. Also, in opposition to the linear scenario, we find a strong dependence on $\Delta b$, particularly when considering a split $\Delta b_{\rm all}$. For $\omega_{\rm cdm}$ and $h$, the degeneracy score is substantially reduced as the bias difference increases, suggesting that this is a key factor in the improved MT constraints relative to ST. The decreasing degeneracy observed for MT indicates that a large bias difference leads to a more orthogonal basis for the cosmological parameters. Interestingly, the score for $A_{s}$ initially increases and then decreases again, despite decreasing error bars (see Fig. 1), indicating that the improvement in $A_{s}$ may be caused by other effects, such as cosmic variance cancellation.
Figure 6: Parameter correlation score defined in Eq. (3.9) as a function of $\Delta b$ for $\omega_{\rm cdm}$ (top), $h$ (middle) and $A_{s}$ (bottom). In the left panels, we consider linear theory, for which we find almost no dependence on $\Delta b$, while in the right panels we use the EFT non-linear modeling. The ST results are shown in black for comparison. The green line indicates a split in all bias parameters $\Delta b_{\rm all}$, while the blue line a simple split in $b_{1}$. Dotted lines show results with the cross-stochasticity removed.
@@ 4.1 More than two tracers
Figure 7: Relative errors in $\omega_{\rm cdm}$, $h$ and $A_{s}$ as a function of the number of tracers $N_{\rm Tracers}$ and for different $\bar{n}^{T}$ values. We fix $\Delta b_{\rm all}=0.6$. For the solid lines, we include the cross-stochastic terms and we exclude them for the dashed lines.
We display as solid lines in Fig. 7 the evolution of the Fisher estimated error in $\omega_{\rm cdm}$, $h$ and $A_{s}$ as a function of the number of tracers, having $\Delta b_{\rm all}=0.6$ fixed. For the tracer split with more tracers, we generalize Eq. (2.27) using evenly spaced MT bias parameters
@@ 4.2 The dependence on the total and relative number density
Figure 8: In the left panel, the relative error in the cosmological parameters as a function of the total number density $\bar{n}^{T}$. Dashed lines correspond to linear theory while solid lines represent the EFT non-linear theory. For the EFT MT analysis, we consider splits in $b_{1}$ and $b_{\rm all}$. In the right panel, the same results normalized by the corresponding ST scenarios. We fix $\Delta b_{1}=0.6$ or $\Delta b_{\rm all}=0.6$.
@@
Figure 9: Error in the cosmological parameters normalized by the ST error as a function of the fraction of the total number density in tracer $t_{1}$ for a non-balanced split, defined as $f_{t_{1}}=\bar{n}^{t_{1}}/\bar{n}^{T}$. Different colors correspond to different values of $\bar{n}^{T}$, in units of $h^{3}/$Mpc3. As before, we fix $\Delta b_{\rm all}=0.6$.
When performing a MT split, we also have the freedom to chose the number densities of the final tracers with respect to the initial sample. In [46, 47] we have only consider a balanced split, in which each subtracer has the same number density. Here we extend this analysis by studying the MT performance as a function of the tracer $t_{1}$ fraction, defined in Eq. (3.2). In this work, we keep the MT relative bias difference $\Delta b$ constant with $f_{t_{1}}$. However, it is important to note that in realistic scenarios where the two samples are split based on a tracer feature, $\Delta b$ is expected to change with $f_{t_{1}}$ (see Sec. 3.2). Since this dependence heavily relies on the sample considered, we assume for simplicity a constant bias difference. We display in Fig. 9 the error as a function of $f_{t_{1}}$ for different values of $\bar{n}^{T}$ normalized by the ST error. Note that the case $f_{t_{1}}\to 0$ or $f_{t_{1}}\to 1$, for which the full density is in one of the samples, does not reproduce the ST case for the same fiducial bias parameters Eq. (2.26), but rather corresponds to a different set of bias parameters shifted by $\Delta b$. We see a typical $U$-shape curve with a relatively flat region for $0.1\lesssim f_{t_{1}}\lesssim 0.9$, which is especially present as we increase $\bar{n}^{T}$. It indicates that, despite the optimal value normally being close to the balanced case $f_{t_{1}}=0.5$, the MT improvements are also present with some unbalancing between both tracer densities. Notice that, in general, it is easier to find subsamples with distinct bias parameters if allowing for their number densities to not be exactly the same. One can, for instance, find the 10 or $20\%$ galaxies with lower or higher values for $b_{\mathcal{G}_{2}}$ and already benefit from the MT information gains.
@@ 5 FoG and different $k_{\rm max}$
Motivated by [100], which treats the zero crossing of the quadrupole a proxy for the perturbative failing of the FoG treatment, we adopt different values of $k_{\rm max}$ for the different samples. We show in Tab. 2 the zero-crossing value for the different samples. We notice that a relatively small shift $\Delta c_{{\rm ct},22}=0.2$ allows for an increase in $k_{\rm max}$ from 0.124h/h/Mpc to 0.208h/h/Mpc. We are assuming that the perturbative scale for dark matter $k_{\rm NL}$ and the higher-derivative expansion scale $1/R_{\rm Halo}$, associated with the Lagrangian halo scale, are smaller than the scales considered here, $k_{\rm NL}>k_{\rm max}$ and $1/R_{\rm Halo}>k_{\rm max}$. We display in the right panel of Fig. 10 the expected error bars for the different $k_{\rm max}$ considered in Tab. 2. We find a notable improvement compared to the ST case due to the tracer split and the higher $k_{\rm max}$ of the bluer sample. This result points out to MT as a natural way to improve the perturbative treatment for FoG.
@@
Table 2: Different $k_{\rm max}$ values (in units of h/h/Mpc) used for the auto, cross and single tracer spectra based on the zero-crossing of the quadrupole.
@@ 6 Forecasts
summing over multiple redshift bins $z_{i}$ and anchoring the power spectra at k=0.1h/k=0.1\,h/Mpc. In addition, we compute the effective number density $\overline{n}_{\text{eff}}$ and the effective redshift $z_{\text{eff}}$ by averaging $\overline{n}(z_{i})$ and $z_{i}$ weighted by the number of tracers $N_{i}=\overline{n}(z_{i})V(z_{i})$ in the corresponding redshift bin. For the linear bias, we use $b_{1}^{\rm eff}(z_{\rm eff})=0.9+0.4z_{\rm eff}$, which has been used in other works [8, 101] and is based on [102]. For simplicity, we fix the other non-linear bias parameters to Eq. (2.26). We summarize in Tab. 3 the surveys and their specifications used in this section.
@@
The forecast results are presented in Fig. 11 as a function of $\Delta b_{\rm all}$. The solid lines represent the most conservative scenario in which we include all the terms, while the dashed lines represent the most optimistic case in which the cross-stochastic component is neglected. We find a factor 2 improvement in $A_{s}$ across all surveys. We also find significant improvement in $h$ and $\omega_{\rm cdm}$, but somewhat smaller for PFS and MegaMapper, which we checked to be due the high $z$ considered for those surveys. The non-linear corrections are very small at high redshift at k=0.15h/k=0.15\,h/Mpc, and typically one can reach higher values of $k_{\rm max}$ at high $z$. A more comprehensive study which considers the different redshift bins of each survey and models their respective $k_{\rm max}$ values as a function of redshift, is left for future work. Finally, when the cross-stochastic component is neglected, we find even stronger MT constraints, highlighting the importance of adding simulation-based priors on these parameters.
@@ 7 Conclusions
In this article, we have extended the results of [46, 47] carrying a systematic study of MT when considering the non-linear modeling of galaxy clustering. We employ a series of Fisher studies to investigate different scenarios, including various survey configurations. Within the large-scale bias expansion, any tracer split based on a specific sample feature can be mapped into a direct split in the bias parameters. Accordingly, we display our results as a function of the difference $\Delta b$ in the bias coefficients and show that the MT Fisher information reproduces the ST scenario when taking $\Delta b\to 0$. Further, we discuss how likely it is to identify two halo samples with differing non-linear bias parameters. While such differences are hard to find in a vanilla scenario, the presence of assembly bias and the use of realistic galaxy samples can lead to $\Delta b_{\mathcal{G}_{2}}\sim 1$ [64], at least when considering samples with large values of $b_{1}$. This paper paves the way toward identifying a MT bias split that does not necessarily rely on finding two samples with very different linear bias coefficients.
@@
Next, we study the importance of cross-stochasticity for MT. While neglecting this term is vital for the gains of MT within linear theory, as pointed out by [75, 20], we show that this is not the case when considering the non-linear bias modeling. Furthermore, we show for the first time, using non-linear modeling, that a split into two tracers seems to maximize the information gain, at least when restricted to a power spectra analysis. We have also found that, when considering the non-linear operators, the MT gains are present at very realistic number density values, in contrast with the findings on linear theory that demands very high tracer densities. We addressed the question of what are the optimal number densities between the two subtracers, finding substantial gains even when considering a very unbalanced split, i.e. when one the tracers is more dense than the other. This result makes the task of finding two tracers with two different set of biases simpler, since we can limit the MT searches to find a small in-homogeneous subsample. Finally, in Sec. 5 we consider the possibility of using different values of $k_{\rm max}$ for the subtracers depending on the strength of the FoG effect for each. We show how this approach can substantially enhance the information extracted from the bluer sample.
We include a summary of the novel results of this work:
  • We present a systematic study of the multi-tracer approach, incorporating full-shape one-loop corrections. While [46, 47] focused on specific simulation setups, e.g. fixed number densities and tracer splits based on halo mass or star formation rate, we explore a wider range of tracer number densities, consider direct splits in the bias parameters and also different noise scenarios.
  • We find, for the first time, that splitting tracers according to their non-linear bias parameters yields results that are comparable to, and often better than, a simple split in $b_{1}$. This extends the standard MT framework, which has primarily focused on selecting samples with distinct linear bias, and opens new possibilities for sample selection strategies aimed at maximizing the information extracted from galaxy clustering. We identify assembly bias in the non-linear coefficients as a promising avenue for identifying samples with differing non-linear bias.
  • We systematically explore the dependence on sample number density for the non-linear case. While, in the linear MT case, improvements only appear at number densities well above $10^{-3}h^{3}$Mpc-3, we find that non-linear MT outperforms ST systematically already at $\bar{n}\sim 10^{-4}h^{3}$Mpc-3.
  • We investigate, for the first time in the context of non-linear MT, the optimal subsample split as a function of number density, going beyond the $50\%/50\%$ balanced split adopted in [46, 47]. We find that even a highly unbalanced split, with $90\%/10\%$ of the total number density, can capture a large fraction of the MT information gain, making it considerably easier to identify distinct subsamples.
  • While the inclusion of cross-stochasticity has traditionally hindered MT performance in the linear full-shape case [75, 20], we find that this is no longer true when non-linear bias terms are included.
  • For the first time, we consider more than two tracers in the non-linear modeling of multi-tracer analyses and find that extending beyond two tracers provides only minor gains.
  • We suggest that a tracer split that leads to different FoG scales, in the lines of [100], can be casted as different $k_{\text{max}}$ choices for subtracers leading to narrower posteriors in cosmological parameters.
  • We forecast the multi-tracer information gain for both current and future survey specifications.
@@ Acknowledgments
We thank Raul Abramo, David Alonso, Yan-Chuan Cai, Anna Cremaschi, Boryana Hadzhiyska, Steffen Hagstotz, Eiichiro Komatsu, Fiona McCarthy, Thiago Mergulhão, Srinivasan Sankarshana, Barbara Sartoris, Fabian Schmidt, Blake Sherwin, Beatriz Tucci and Matteo Zennaro for useful discussions. We thank Alex Barreira, Mathias Garny and Rodrigo Voivodic for valuable comments in the draft.
@@ Appendix A Stability of Fisher derivatives
Figure 12: Relative error $\sigma_{h}/h^{\rm fid}$, $\sigma_{\omega_{\rm cdm}}/\omega_{\rm cdm}^{\rm fid}$ and $\sigma_{A_{s}}/A_{s}^{\rm fid}$ as a function of the relative step size $\delta h$, $\delta\omega_{\rm cdm}$ and $\delta A_{s}$ used in the (second-order) finite difference method. A small step size is consistent with noise while a large step size leads to the breakdown of the finite difference method. We adopt $\delta h/h^{\rm fid}=\delta\omega_{\rm cdm}/\omega_{\rm cdm}^{\rm fid}=\delta A_{s}/A_{s}^{\rm fid}=0.04$, where the errors remain relatively stable.
The derivatives $\partial_{\theta_{a}}P^{t_{i}t_{j}}(k_{\alpha})$ in the Fisher analysis of Eq. (2.18) can be computed analytically with respect to the bias and stochastic parameters but have to be computed numerically for $h$, $\omega_{\rm cdm}$ and $A_{s}$. To improve the stability of the numerical derivatives, we use a second-order finite difference method and compute the log-derivative $P^{t_{i}t_{j}}(k_{\alpha})\partial_{\theta_{a}}\ln P^{t_{i}t_{j}}(k_{\alpha})$. We show in Fig. 12 the stability of the Fisher-extracted relative parameters errors $\sigma_{h}$, $\sigma_{\omega_{\rm cdm}}$ and $\sigma_{A_{s}}$ as a function of the step sizes $\delta h$, $\delta\omega_{\rm cdm}$ and $\delta A_{s}$ used for the finite difference evaluation. We observe that choosing a step size that is too small results in numerical noise, while large step sizes lead to the breakdown of the finite difference method. Intermediate values lead to relatively stable results. Based on this, we adopt the relative step size $\delta h/h^{\rm fid}=\delta\omega_{\rm cdm}/\omega_{\rm cdm}^{\rm fid}=\delta A_{s}/A_{s}^{\rm fid}=0.04$, normalized by their fiducial Planck 18 values.